OpenStudy (anonymous):

Please solve quickly and completely!!! I'LL FAN U AND MEDAL U!!!!!! ALGEBRA 1 Part A: Solve C = 5/9 (F - 32) for F. Part B: Determine the value of F when C = 8 °C. Part C: Solve –np – 80 > 60 for n. Show your work.

OpenStudy (studygurl14):

\(\large C=\Large\frac{5}{9}(F-32)\) First step, multiply both sides by 9

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

Must sound pretty dumb, but I'm really bad at math...

OpenStudy (studygurl14):

Just multiply both sides by 9...I don't know how else to explain it... \(\large 9C=(9)\Large\frac{5}{9}(F-32)\) the 9s cancel out on the right side \(\large 9C=\Large 5(F-32)\)

OpenStudy (studygurl14):

similarily, now divide both sides by 5

OpenStudy (usukidoll):

why not multiply 5 (F-32) to make it 9C = 5F- 160? Then add 160 on both sides and divide by 5

OpenStudy (usukidoll):

you need F by itself so the final answer should be whatever it is = F

OpenStudy (studygurl14):

that would work too. doesn't really matter what way you do it

OpenStudy (studygurl14):

my way has less steps though

OpenStudy (anonymous):

Ok!!!

OpenStudy (anonymous):

Thanks for your awesomeness!!!

OpenStudy (anonymous):

Somehow the one with the most steps helped me more... But thanks anyway!!! ^-^

OpenStudy (usukidoll):

:O thanks

OpenStudy (anonymous):

No prob!

OpenStudy (studygurl14):

Nice job then @UsukiDoll :) lol